OpenStudy (anonymous):

HELP i rlly dont get it the volume of a pyramid with a rectangular base is v=1/3LWH. find W if V=500 L=15 and H=4. a) W=25 b) W=20 c) W=18 d) W=15

OpenStudy (anonymous):

500 = 1/3 x (15)(4)(W) 500 = 1/3 x 60W 1500 = 60W W = 25

OpenStudy (anonymous):

thank you.